OpenStudy (anonymous):

could you also help me with (-1)^(n+1)(11^(n-1))/((n+7)^6(10^(n+2)) for L and where the equation editor is?

OpenStudy (loser66):

I don't get what you mean by where the equation editor is? Your problem : \[L=\dfrac{(-1)^{n+1}*(11^{n-1})}{(n+7)^6*(10^{n+2})}\] right? and what are we supposed to do?

OpenStudy (anonymous):

find L value

OpenStudy (anonymous):

thanks!

OpenStudy (loser66):

find L value for n =??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

looking to find the value of L in the ratio test

OpenStudy (loser66):

You meant test whether the sequence converges or diverges?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

Not necessarily related to your problem, but you asked where the equation editor is, you'll find it here: (see attached)

OpenStudy (loser66):

@jtvatsim please help him, I forgot how to do ratio test. :)

OpenStudy (jtvatsim):

Ah... I was hoping you knew... Let me dust off the Calculus part of my brain. If I recall, it's pretty straightforward. :)

OpenStudy (jtvatsim):

The ratio test says that we evaluate \[L = \lim_{n \rightarrow \infty} | \frac{a_{n+1}}{a_n} |\]

OpenStudy (jtvatsim):

The criteria are that if L < 1 the series converges absolutely, if L > 1 the series diverges, and if L = 1, we have no clue.

OpenStudy (jtvatsim):

So, with your current question, we have a very ugly fraction to take care of, but let's see what we get...

OpenStudy (jtvatsim):

I'm not going to type out the gigantic intermediate steps, but I will clarify if you get an answer different than mine.

OpenStudy (jtvatsim):

OK, I've gotten to this point so far

OpenStudy (jtvatsim):

\[\lim_{n \rightarrow \infty} | - \frac{11 (n+7)^6}{10 (n+8)^6} |\]

OpenStudy (jtvatsim):

Is that somewhat similar to what you got after simplifying?

OpenStudy (anonymous):

yes you don't have to do the steps just struggling with may answer part, i have gotten many problems right so far but for some reason can't get this one

OpenStudy (anonymous):

any updates???